# Bayesian Prompt Flow Learning for Zero-Shot Anomaly Detection

### v1

@@ Abstract
Recently, vision-language models (e.g. CLIP) have demonstrated remarkable performance in zero-shot anomaly detection (ZSAD). By leveraging auxiliary data during training, these models can directly perform cross-category anomaly detection on target datasets, such as detecting defects on industrial product surfaces or identifying tumors in organ tissues. Existing approaches typically construct text prompts through either manual design or the optimization of learnable prompt vectors. However, these methods face several challenges: 1) handcrafted prompts require extensive expert knowledge and trial-and-error; 2) single-form learnable prompts struggle to capture complex anomaly semantics; and 3) an unconstrained prompt space limit generalization to unseen categories. To address these issues, we propose Bayesian Prompt Flow Learning (Bayes-PFL), which models the prompt space as a learnable probability distribution from a Bayesian perspective. Specifically, a prompt flow module is designed to learn both image-specific and image-agnostic distributions, which are jointly utilized to regularize the text prompt space and enhance the model’s generalization on unseen categories. These learned distributions are then sampled to generate diverse text prompts, effectively covering the prompt space. Additionally, a residual cross-attention (RCA) module is introduced to better align dynamic text embeddings with fine-grained image features. Extensive experiments on 15 industrial and medical datasets demonstrate our method’s superior performance.
@@ 1 Introduction
In the fields of industrial defect detection [39, 4, 18] and medical image analysis [2, 11, 28], the cold-start problem is a major challenge, characterized by insufficient labeled data for supervised training on new object categories. An effective solution is zero-shot anomaly detection (ZSAD), which directly detect anomalies on any new object surface after learning from auxiliary data. However, significant variations in background features, anomaly types, and visual appearances across different products and organs make it difficult to achieve robust generalization.
Vision-language models (VLMs), such as CLIP [32] and ALIGN [20], trained on large-scale image-text pairs through contrastive representation learning, have shown significant potential for ZSAD. As a representative VLM, CLIP [32] has been widely adopted in previous works [18, 46, 8]. A key challenge for CLIP-based ZSAD methods is the effective design of text prompts. As illustrated in Figure 1(a), a typical textual prompt consists of three components: context words, state words, and class. They respectively represent a general description of the object (e.g. background), the attributes indicating whether it is normal or anomalous (e.g. good / damaged), and the object category (e.g. wood). Existing text prompt design methods can be broadly categorized into two types: prompt ensemble-based and prompt optimization-based. Prompt ensemble-based methods, such as WinCLIP [18], APRIL-GAN [9], and CLIP-AD [10], create handcrafted prompt templates and embeds various context and state words into these templates to generate a large number of textual prompts. However, the number of context and state words is limited and the design relies on expert knowledge. In contrast, prompt optimization-based methods, such as AnomalyCLIP [46] and AdaCLIP [8], replace the context words in the prompts with learnable vectors or directly insert them into the encoder. However, two critical problems are often overlooked: 1) The design of learnable prompts is overly simplistic, which makes it difficult to capture complex context semantics; 2) The learnable prompt space is not properly constrained, making it challenging to generalize learnable text prompts to unseen categories during training.
@@
First, we construct two-class prompt banks to store learnable textual prompts for normal and abnormal cases. Then, a prompt flow module is designed to learn the distribution of context and state words within the text prompts. Specifically, to introduce rich visual semantics into the text prompts, ISD is learned through prompt flow to model the context words, dynamically adapting based on the input image. To learn unified normal and abnormal semantics, IAD is established for the state words, remaining static and conditioned on a learnable free vector. Finally, multiple textual prompts sampled from the learned distribution are fused with the prompt banks to generate enriched text prompts. Moreover, to align the dynamically sampled text embeddings with fine-grained image features, we introduce a residual cross-modal attention (RCA) module to enhance the interaction between modalities. Unlike existing methods [9, 10, 8] that directly align text with patch-level features as shown in Figure 1(a), our RCA module updates dynamic text embeddings with fine-grained image features to improve the ZSAD performance. The results generated from alignment of multiple updated text embeddings and patch-level feature are integrated to produce the final anomaly map.
The contributions of this work are summarized as follows: 1) We propose a novel Bayesian Prompt Flow Learning method, Bayes-PFL, built upon CLIP to address the ZSAD problem. By distribution sampling and distribution regularization of the text prompt space, Bayes-PFL improves the model’s ZSAD performance on novel categories; 2) The prompt flow module is designed to learn ISD and IAD, which model the context and state semantics in text prompts, respectively. Additionally, the RCA module is introduced to improve the alignment between fine-grained features and dynamic text prompts through cross-modal interactions; 3) After training on auxiliary datasets, our method can directly detect defects / lesions on novel products / organ without requiring target training data. Extensive experiments across 15 industrial and medical datasets demonstrate that Bayes-PFL achieves state-of-the-art (SOTA) performance in ZSAD.
@@ 2.1 Zero-shot Anomaly Detection
The earliest work, WinCLIP [18] proposes a window-based strategy to aggregate classification results obtained from aligning text with sub-images at different scales. APRIL-GAN [9] and CLIP-AD [10] use trainable adapter layers to map fine-grained patch features into a joint embedding space. The latter further employs feature surgery to address the issue of opposite predictions. Additionally, several existing studies have integrated multiple foundational models, such as SAM [24] and GroundingDINO [27], to collectively enhance ZSAD performance, as demonstrated in SAA/SAA+ [7] and ClipSAM [26]. Recently, an increasing number of prompt optimization-based methods have been proposed. AnomalyCLIP [46] and Filo [13] insert learnable vectors into the input text or CLIP encoder layers to avoid extensive engineering on hand-crafted prompt design. AdaCLIP [8] and VCP-CLIP [31] further utilize textual and visual hybrid prompts to enhance CLIP’s anomaly perception capability. Different from these methods, our Bayes-PFL learns the distribution of textual prompts and leverages sampling to cover the prompt space to improve the ZSAD performance.
@@ 2.2 Prompt Design
The design of textual prompts plays a pivotal role in applying VLMs to downstream tasks. Manually crafted textual prompts [32, 15, 40, 25, 42] and data-driven approaches to obtain learnable prompts [45, 44, 33, 41, 14] in the natural scenario have both demonstrated effectiveness in zero-shot learning. In the ZSAD task, Compositional Prompt Ensemble [18, 9, 10] and Prompt Optimization [46, 31, 8] are two typical methods for text prompt design.
Compositional Prompt Ensemble generates diverse textual prompts through the ensemble of state words and text templates. In the representative work APRIL-GAN [9], approximately 35 templates and 5 state words are designed for abnormal descriptions, resulting in a total of $35\times 5$ textual prompts. The text embeddings from these prompts are averaged to produce the final result, and a similar approach is applied to normal descriptions.
Prompt Optimization replaces the context words in text prompts with learnable vectors or directly inserting these vectors into the encoder layers. As depicted in Figure 1(a), the representative method AnomalyCLIP [46] designs text prompts in the form of $[\mathbf{V}_{1}][\mathbf{V}_{2}]\cdots[\mathbf{V}_{l}][\text{state}][\text{object}]$, where $\mathbf{V}_{i},i=1,2,\cdots,l$ are learnable vectors. However, the single-form prompt design and the lack of effective constraints on the prompt space limit its generalization capability for unseen categories.
@@ 3.1 Overview
Figure 2 illustrates the main framework of Bayes-PFL, which consists of three key designs: 1) Two-class prompt banks for normal and abnormal textual descriptions, 2) A prompt flow module for distribution learning, and 3) A RCA module for feature alignment. Given an input image $\mathbf{X}\in\mathbb{R}^{h\times w\times 3}$, the model first map the patch-level image features to the joint embedding space through a single linear layer, producing $\mathbf{F}^{I}\in\mathbb{R}^{H\times W\times C}$, where $H=h/patchsize,W=w/patchsize$ and $C$ represents the dimensionality of the joint embedding space. Subsequently, the prompt flow module learns the image-specific distribution $q_{K}(\boldsymbol{\varphi}^{e})$ for context words, and the image-agnostic distributions $q_{K}(\boldsymbol{\varphi}^{n}),q_{K}(\boldsymbol{\varphi}^{a})$ for normal and abnormal state words. Monte Carlo sampling [1] is applied to these distributions, and the sampled results $\boldsymbol{\varphi}^{e}_{r},\boldsymbol{\varphi}^{n}_{r},\boldsymbol{\varphi}^{a}_{r}$ are fused with the prompt banks to generate diverse textual prompts. The text embeddings $\mathbf{Z}_{b,r}^{t}$ are aligned with $\mathbf{F}^{I}$ to obtain multiple anomaly segmentation results, which are averaged to acquire the final anomaly map $\mathbf{M}\in\mathbb{R}^{H\times W}$. Note that when aligning with $\mathbf{F}^{I}$, we additionally employ the designed RCA module to refine the text embeddings to $\mathbf{F}_{b,r}^{t}$.
@@ 3.2 Two-class Prompt Banks
We decompose the text prompt into three components: context words, state words, and class words. Unlike previous prompt optimization strategies [46] that manually design state words and only optimize context words, we argue that state words (e.g. good / damaged) still reside within an optimizable prompt space. Thus, in the word embedding space, the manually designed context and state words are replaced with learnable vectors. Inspired by Compositional Prompt Ensemble, normal and abnormal prompt banks are further constructed to expand the prompt space, with each containing $B$ distinct trainable prompts:
@@
where $b=1,2,\cdots,B$. $\mathbf{E}_{b,i}\in\mathbb{R}^{C},i=1,2,\cdots,P$ are learnable vectors designed to encode contextual information from the text. $\mathbf{S}_{m,j}^{n},\mathbf{S}_{m,j}^{a}\in\mathbb{R}^{C},j=1,2,\cdots,Q$ are learnable normal and abnormal state vectors, respectively. Note that we enforce orthogonality among the text embeddings generated from the same prompt bank, promoting diversity in the normal or anomalous semantics captured by different prompts. This constraint is implemented through the orthogonal loss $\mathcal{L}_{ort}$ introduced in Section 3.6.
@@ 3.3 Prompt Distributionalization
Bayesian inference. Assume the auxiliary training dataset is $D=\{\mathbf{X},Y_{c},\mathbf{Y}_{s}\}$, where $\mathbf{X}\in\mathbb{R}^{h\times w\times 3}$ represents the input image, and $Y_{c}\in\{0,1\}$, $\mathbf{Y}_{s}\in\mathbb{R}^{h\times w}$ denote the image-level labels and pixel-level ground truth, respectively. To construct the text prompt distribution, the context, normal state, and abnormal state word embeddings are represented as $C$-dimensional random vectors $\Phi_{e}$, $\Phi_{s}^{n}$ and $\Phi_{s}^{a}$, respectively. For convenience, let $\Phi=\{\Phi_{e},\Phi_{s}^{n},\Phi_{s}^{a}\}$. The posterior probabilities are then computed as:
@@
Since calculating $p(D)$ is tractable, variational inference [21] with distribution $q_{\gamma}(\Phi)$, parameterized by $\gamma$, is employed to approximate the posterior probability. Note that our method operates under the mean-field assumption, where all variables are mutually independent: $q_{\gamma}(\Phi)=q_{\gamma}(\boldsymbol{\varphi}^{e})q_{\gamma}(\boldsymbol{\varphi}^{n})q_{\gamma}(\boldsymbol{\varphi}^{a})$. By applying Jensen’s inequality, an upper bound is drived for the log marginal likelihood of the training data:
@@
The variational distribution $q_{\gamma}(\Phi)$ can be acquired by minimizing $\mathcal{L}_{e}(D)$.
Prompt flow module. Richer posterior approximations enable us to better estimate the prompt distribution [34]. Hence, a prompt flow module is designed to transform a simple probability distribution $q_{0}(\Phi_{0})$ into a more complex one $q_{K}(\Phi_{K})$ through a series of invertible mappings $h_{k}$: $\Phi_{K}=h_{K}(h_{K-1}(\cdots h_{1}(\Phi_{0})))$, as shown in the bottom right corner of Figure 2. For the model efficiency, we use a linear-time transformation $h(\Phi)=\Phi+\mathbf{u}g(\mathbf{w}^{\text{T}}\Phi+\mathbf{b})$, where $\mathbf{w},\mathbf{u}\in\mathbb{R}^{C}$, $\mathbf{b}\in\mathbb{R}$ are trainable parameters and $g(\cdot)$ is the Tanh activation function [36]. The new distribution after $K$ transformations is expressed as:
@@
where the prior follows $p=\mathcal{N}(0,\mathbf{I})$. The initial density $q_{0}$ is defined as a standard normal distribution: $q_{0}=\mathcal{N}(\boldsymbol{\mu}(\boldsymbol{\xi}),\boldsymbol{\Sigma}(\boldsymbol{\xi}))$, where $\boldsymbol{\mu}\in\mathbb{R}^{C}$ and $diag(\boldsymbol{\Sigma})=\boldsymbol{\sigma}\in\mathbb{R}^{C}$ are conditioned to the vector $\boldsymbol{\xi}\in\mathbb{R}^{C}$ and parameterized by three linear layers, with a Softplus activation function [36] applied between consecutive layers.
@@
Prompt sampling and fusion. Monte Carlo sampling is used to sample $R$ iterations from the initial densities $q_{0}^{e},q_{0}^{n},q_{0}^{a}$. The sampled results are processed through the prompt flow module to obtain $\boldsymbol{\varphi}^{e}_{r},\boldsymbol{\varphi}^{n}_{r},\boldsymbol{\varphi}^{a}_{r},r=1,2,\cdots,R$, which represent sampled context vector, normal sampled state vector and abnormal sampled state vector, respectively. The fusion process can be shown as:

### v2

@@ Abstract
Recently, vision-language models (e.g. CLIP) have demonstrated remarkable performance in zero-shot anomaly detection (ZSAD). By leveraging auxiliary data during training, these models can directly perform cross-category anomaly detection on target datasets, such as detecting defects on industrial product surfaces or identifying tumors in organ tissues. Existing approaches typically construct text prompts through either manual design or the optimization of learnable prompt vectors. However, these methods face several challenges: 1) handcrafted prompts require extensive expert knowledge and trial-and-error; 2) single-form learnable prompts struggle to capture complex anomaly semantics; and 3) an unconstrained prompt space limits generalization to unseen categories. To address these issues, we propose Bayesian Prompt Flow Learning (Bayes-PFL), which models the prompt space as a learnable probability distribution from a Bayesian perspective. Specifically, a prompt flow module is designed to learn both image-specific and image-agnostic distributions, which are jointly utilized to regularize the text prompt space and improve the model’s generalization on unseen categories. These learned distributions are then sampled to generate diverse text prompts, effectively covering the prompt space. Additionally, a residual cross-model attention (RCA) module is introduced to better align dynamic text embeddings with fine-grained image features. Extensive experiments on 15 industrial and medical datasets demonstrate our method’s superior performance. The code is available at https://github.com/xiaozhen228/Bayes-PFL.
@@ 1 Introduction
In the fields of industrial defect detection [41, 42, 18, 32] and medical image analysis [2, 11, 29, 8], the cold-start problem is a major challenge, characterized by an insufficient amount of labeled data for supervised training on new object categories. An effective solution is zero-shot anomaly detection (ZSAD), which directly detects anomalies on new object surface after being trained on auxiliary data. However, significant variations in background features, anomaly types, and visual appearances across different products and organs make it difficult to achieve robust generalization.
Vision-language models (VLMs), such as CLIP [34] and ALIGN [20], which are trained on large-scale image-text pairs through contrastive representation learning, have shown significant potential for ZSAD. As a representative VLM, CLIP [34] has been widely adopted in previous works [18, 49, 8]. A key challenge for CLIP-based ZSAD methods is the effective design of text prompts. As illustrated in Figure 1(a), a typical textual prompt consists of three components: context words, state words, and class. They respectively represent a general description of the object (e.g. background), the attributes indicating whether it is normal or anomalous (e.g. good / damaged), and the object category (e.g. wood). Existing text prompt design methods can be broadly categorized into two types: prompt ensemble-based and prompt optimization-based. Prompt ensemble-based methods, such as WinCLIP [18], APRIL-GAN [9], and CLIP-AD [10], create handcrafted prompt templates and embed various context and state words into these templates to generate a large number of textual prompts. However, the number of context and state words is limited and the design relies on expert knowledge. In contrast, prompt optimization-based methods, such as AnomalyCLIP [49] and AdaCLIP [8], replace the context words in the prompts with learnable vectors or directly insert them into the encoder. However, two critical issues are often overlooked: 1) The design of learnable prompts is overly simplistic, which makes it difficult to capture complex context semantics; 2) The learnable prompt space is not properly constrained, hindering the generalization of learnable text prompts to unseen categories.
@@
First, we construct two-class prompt banks to store learnable textual prompts for normal and abnormal cases. Then, a prompt flow module is designed to learn the distribution of context and state words within the text prompts. Specifically, to introduce rich visual semantics into the text prompts, ISD is learned through prompt flow to model the context words, dynamically adapting based on the input image. To encode unified normal and abnormal semantics, IAD is established for the state words, remaining static and conditioned on a learnable free vector. Finally, multiple textual prompts sampled from the learned distribution are fused with the prompt banks to generate enriched text prompts. Moreover, to align the sampled text embeddings with fine-grained image features, we introduce a residual cross-modal attention (RCA) module to enhance the interaction between modalities. Unlike existing methods [9, 10, 8] that directly align text with patch-level features as shown in Figure 1(a), our RCA module updates dynamic text embeddings using fine-grained image information to improve the ZSAD performance. The results generated from alignment of multiple updated text embeddings and patch-level feature are integrated to produce the final anomaly map.
The contributions of this work are summarized as follows: 1) We propose a novel Bayesian Prompt Flow Learning method, Bayes-PFL, built on CLIP to address the ZSAD problem. By distribution sampling and distribution regularization of the text prompt space, Bayes-PFL enhances the ZSAD performance on novel categories; 2) The prompt flow module is designed to learn ISD and IAD, which model the context and state semantics in text prompts, respectively. Additionally, the RCA module is introduced to improve the alignment between fine-grained features and dynamic text prompts through cross-modal interactions; 3) After training on auxiliary datasets, our method can directly detect defects / lesions on novel products / organs without requiring target training data. Extensive experiments across 15 industrial and medical datasets demonstrate that Bayes-PFL achieves state-of-the-art (SOTA) performance in ZSAD.
@@ 2.1 Zero-shot Anomaly Detection
The earliest work, WinCLIP [18], proposes a window-based strategy to aggregate classification results obtained from aligning text with sub-images at multiple scales. APRIL-GAN [9] and CLIP-AD [10] use trainable adapter layers to map fine-grained patch features into a joint embedding space. The latter further employs feature surgery to mitigate opposite predictions. Additionally, several existing studies have integrated multiple foundation models, such as SAM [25] and GroundingDINO [28], to collectively enhance ZSAD performance, as demonstrated in SAA/SAA+ [7] and ClipSAM [27]. Recently, an increasing number of prompt optimization-based methods have been proposed. AnomalyCLIP [49] and Filo [13] insert learnable vectors into the input text or CLIP encoder layers to avoid extensive engineering on hand-crafted prompt design. AdaCLIP [8] and VCP-CLIP [33] further utilize textual and visual hybrid prompts to enhance CLIP’s anomaly perception capability. Different from these methods, our Bayes-PFL learns the distribution of textual prompts and leverages sampling to cover the prompt space to improve the ZSAD performance.
@@ 2.2 Prompt Design
The design of textual prompts plays a pivotal role in applying VLMs to downstream tasks. Manually crafted textual prompts [34, 15, 43, 26, 45] and data-driven approaches to obtain learnable prompts [48, 47, 35, 44, 14] in the natural scenario have both demonstrated effectiveness in zero-shot learning. In the ZSAD task, Compositional Prompt Ensemble [18, 9, 10] and Prompt Optimization [49, 33, 8] are two mainstream methods for text prompt design.
Compositional Prompt Ensemble generates diverse text prompts through the combination of state words and textual templates. In the representative work APRIL-GAN [9], approximately 35 templates and 5 state words are designed for abnormal descriptions, resulting in a total of $35\times 5$ text prompts. The text embeddings from these prompts are averaged to produce the final result, and a similar approach is applied to normal descriptions.
Prompt Optimization replaces the context words in text prompts with learnable vectors or directly inserts these vectors into the encoder layers. As depicted in Figure 1(a), the representative method AnomalyCLIP [49] designs text prompts in the form of $[\mathbf{V}_{1}][\mathbf{V}_{2}]\cdots[\mathbf{V}_{z}][\text{state}][\text{object}]$, where $\mathbf{V}_{i},i=1,2,\cdots,z$ are learnable vectors. However, the single-form prompt design and the lack of effective constraints on the prompt space limit its generalization capability to unseen categories.
@@ 3.1 Overview
Figure 2 illustrates the main framework of Bayes-PFL, which consists of three key designs: 1) Two-class prompt banks for normal and abnormal textual descriptions, 2) A prompt flow module for distribution learning, and 3) A RCA module for feature alignment. Given an input image $\mathbf{X}\in\mathbb{R}^{h\times w\times 3}$, the model first maps the patch-level image features to the joint embedding space through a single linear layer, producing $\mathbf{F}^{I}\in\mathbb{R}^{HW\times C}$, where $H=h/patchsize,W=w/patchsize$ and $C$ represents the dimensionality of the joint embedding space. Subsequently, the prompt flow module learns the image-specific distribution $q_{K}(\boldsymbol{\varphi}^{e})$ for context words, and the image-agnostic distributions $q_{K}(\boldsymbol{\varphi}^{n}),q_{K}(\boldsymbol{\varphi}^{a})$ for normal and abnormal state words. Monte Carlo sampling [1] is applied to these distributions, and the sampled results $\boldsymbol{\varphi}^{e}_{r},\boldsymbol{\varphi}^{n}_{r},\boldsymbol{\varphi}^{a}_{r}$ are fused with the prompt banks to generate diverse textual prompts. The text embeddings $\mathbf{Z}_{b,r}^{t}$ are aligned with $\mathbf{F}^{I}$ to obtain multiple anomaly segmentation results, which are averaged to derive the final anomaly map $\mathbf{M}$. Note that when aligning with $\mathbf{F}^{I}$, we additionally employ the proposed RCA module to refine the text embeddings to $\mathbf{F}_{b,r}^{t}$.
@@ 3.2 Two-class Prompt Banks
We decompose the text prompt into three components: context words, state words, and class words. Unlike previous prompt optimization strategies [49], which manually design state words and only optimize context words, we argue that state words (e.g., good / damaged) still reside within an optimizable prompt space. Thus, in the word embedding space, the manually designed context and state words are replaced with learnable vectors. Inspired by Compositional Prompt Ensemble, normal and abnormal prompt banks are further constructed to expand the prompt space, with each containing $B$ distinct trainable prompts:
@@
where $b=1,2,\cdots,B$. $\mathbf{E}_{b,i}\in\mathbb{R}^{C},i=1,2,\cdots,P$ are learnable vectors designed to encode contextual information in the text. $\mathbf{S}_{b,j}^{n},\mathbf{S}_{b,j}^{a}\in\mathbb{R}^{C},j=1,2,\cdots,Q$ are learnable normal and abnormal state vectors, respectively. Note that we enforce orthogonality among the text embeddings generated from the same prompt bank, promoting diversity in the normal or anomalous semantics captured by different prompts. This constraint is implemented through the orthogonal loss $\mathcal{L}_{ort}$ introduced in Section 3.6.
@@ 3.3 Prompt Distributionalization
Bayesian inference. Assume the auxiliary training dataset is $D=\{\mathbf{X},Y_{c},\mathbf{Y}_{s}\}$, where $\mathbf{X}\in\mathbb{R}^{h\times w\times 3}$ represents the input image, and $Y_{c}\in\{0,1\}$, $\mathbf{Y}_{s}\in\mathbb{R}^{h\times w}$ denote the image-level labels and pixel-level ground truth, respectively. To construct the text prompt distribution, the context, normal state, and abnormal state word embeddings are represented as $C$-dimensional random vectors $\Phi_{e}$, $\Phi_{s}^{n}$ and $\Phi_{s}^{a}$, respectively. For convenience, we denote $\Phi=\{\Phi_{e},\Phi_{s}^{n},\Phi_{s}^{a}\}$. The posterior probabilities are then computed as:
@@
Since computing the marginal likelihood $p(D)$ is intractable, variational inference [21] with distribution $q_{\gamma}(\Phi)$, parameterized by $\gamma$, is employed to approximate the posterior probability. Note that our method operates under the mean-field assumption, where all variables are mutually independent: $q_{\gamma}(\Phi)=q_{\gamma}(\boldsymbol{\varphi}^{e})q_{\gamma}(\boldsymbol{\varphi}^{n})q_{\gamma}(\boldsymbol{\varphi}^{a})$. By applying Jensen’s inequality, an upper bound is drived for the log marginal likelihood of the training data:
@@
Thus, the variational distribution $q_{\gamma}(\Phi)$ can be obtained by minimizing the evidence lower bound (ELBO) loss $\mathcal{L}_{e}(D)$.
Prompt flow module. Richer posterior approximations enable us to better estimate the prompt distribution [36]. Hence, a prompt flow module is designed to transform a simple probability distribution $q_{0}(\Phi_{0})$ into a more complex one $q_{K}(\Phi_{K})$ through a series of invertible mappings $h_{k},k=1,2,\cdots K$. Specifically, this transformation follows $\Phi_{K}=h_{K}(h_{K-1}(\cdots h_{1}(\Phi_{0})))$, as illustrated in the bottom right corner of Figure 2. For the model efficiency, we use a linear-time transformation $h(\Phi)=\Phi+\mathbf{u}g(\mathbf{w}^{\text{T}}\Phi+\mathbf{b})$, where $\mathbf{w},\mathbf{u}\in\mathbb{R}^{C}$, $\mathbf{b}\in\mathbb{R}$ are trainable parameters and $g(\cdot)$ is the Tanh activation function [38]. After $K$ transformations, the final distribution is given by:
@@
where the prior distribution follows $p=\mathcal{N}(0,\mathbf{I})$. The initial density $q_{0}$ is defined as a standard normal distribution: $q_{0}=\mathcal{N}(\boldsymbol{\mu}(\boldsymbol{\xi}),\boldsymbol{\Sigma}(\boldsymbol{\xi}))$, where $\boldsymbol{\mu}\in\mathbb{R}^{C}$ and $diag(\boldsymbol{\Sigma})=\boldsymbol{\sigma}\in\mathbb{R}^{C}$ are conditioned on the vector $\boldsymbol{\xi}\in\mathbb{R}^{C}$ and parameterized by three linear layers, with a Softplus activation function [38] applied between consecutive layers.
@@
Prompt sampling and fusion. Monte Carlo sampling is used to sample $R$ iterations from the initial densities $q_{0}^{e},q_{0}^{n},q_{0}^{a}$. The sampled results are processed through the prompt flow module to obtain $\boldsymbol{\varphi}^{e}_{r},\boldsymbol{\varphi}^{n}_{r},\boldsymbol{\varphi}^{a}_{r},r=1,2,\cdots,R$, which represent sampled context vector, normal sampled state vector and abnormal sampled state vector, respectively. The fusion process is formulated as: